If A_1,A_2 and A_3 are independent events,then check whether or not
(i)A_1 and A^c_2 are independent;
(ii)A_1 and A^c_2UA^c_3 are independent..
1
Expert's answer
2019-03-27T11:14:44-0400

(i) Recall that two events "A_i" and "A_j" are independent iff. "P\left(A_i\cap A_j\right)=P\left(A_i\right)P(A_j)". We may check this:

"P\left(A_1\cap {A_2}^c\right)=P\left(A_1\cap \left(A_2\sqcup {A_2}^c\right)\right)-P\left(A_1\cap A_2\right)
=P\left(A_1\right)-P\left(A_1\right)P\left(A_2\right)=P\left(A_1\right)\left(1-P\left(A_2\right)\right)
=P\left(A_1\right)\left(P\left(A_2\sqcup {A_2}^c\right)-P\left(A_2\right)\right)
=P\left(A_1\right)P\left({A_2}^c\right)."


(ii) We are given that set "{\left\{A_k\right\}}_{1,2,3}" is not just pairwise independent, it is mutually independent. Hence "A_1" and "A_2\cap A_3" are independent. From (i), it is equivalent that "A_1" and "{A_2}^c\cup {A_3}^c" are independent too.


Answer:

(i) Yes (if "{\left\{A_k\right\}}_{1,2,3}" is mutually or pairwise independent).

(ii) Yes (since "{\left\{A_k\right\}}_{1,2,3}" is mutually independent).


References:

1 "\url{https://en.wikipedia.org/wiki/Independence_(probability_theory)#For_events}"

2 "\url{https://en.wikipedia.org/wiki/Independence_(probability_theory)#More_than_two_events}"